# DRS: Deep Question Reformulation With Structured Output

### v3

@@ 1 Introduction
We conduct extensive experiments on six diverse datasets, comparing DRS with multiple baseline approaches across different types of questions and domains. The results demonstrate that DRS significantly improves reformulation accuracy across all tested LLMs, with GPT-3.5’s performance increasing from 23.03% to 70.42% and open-source model Gemma2-9B improving from 26.35% to 56.75%. Additionally, we introduce a more reliable evaluation framework using GPT-4o-mini, replacing the previous Llama2-7B evaluator to ensure more accurate assessment of reformulation quality.
@@ 3.1 Entity Extraction and Filtering
Therefore, this process involves two sub-steps:
@@
Table 1: The detail information of all experimental data for evaluating the performance on the question reformulation task.
@@ 3.2 DFS Combination Search and Structured Question Generation
In this step, our algorithm uses a DFS search approach to identify possible entity combinations that can be reformulated into new questions. During the generation process, we find that directly prompting the large language model to reshape the question often results in a low success rate due to the model’s potentially unlimited output content. To address this, we guide the large language model to generate structured outputs by restricting the entities required in the generated text. We ensure the reliability of the reformulated question by first generating a statement, then using that statement to create the question. Additionally, to enhance the algorithm’s efficiency, we verify the entities immediately after generating the new question, preventing unnecessary subsequent steps. We also apply pruning to the DFS algorithm by limiting the total number of preselected questions and controlling the search depth.
@@ 4.1 Datasets
In this paper, we utilize the newly constructed high-quality dataset CouldAsk [^1] https://huggingface.co/datasets/wentingzhao/couldask, introduced by Zhao et al. (2024). This dataset comprises subsets from six diverse sources, including Yelp, BBC, SQuAD, Reddit, and others, featuring over 4,000 high-difficulty, high-quality data points filtered by both human annotators and GPT-4. For our study, we focus on over 1,700 unanswerable questions selected from all subsets as test data for model reformulation. The detailed information about our experimental data is provided in Table 1.
@@ 4.2 Large Language Models
To evaluate our methodology and compare it with baseline approaches from previous studies, we conduct all experiments using four different large language models.
Closed-Source Models. To balance experimental cost and effectiveness, we primarily use two widely adopted models from OpenAI: GPT-3.5 and GPT-4o-mini.
Open-Source Models. We also perform experiments using two well-known models developed by Google and Alibaba: Gemma2-9B Team et al. (2024) and Qwen2.5-7B Yang et al. (2024); Team (2024).
@@ 4.3 Metric
(i) We use GPT-4o-mini as an evaluator to generate reasoning steps and assess whether the reformulated question can be answered using the relevant text. (ii) We also use GPT-4o-mini as an entity detector to measure the overlap of entities between the reformulated question and the original question. If the number of overlapping entities exceeds half of the labeled entities in the dataset, we consider the reformulated question likely to help the inquirer obtain the desired information.
@@
4.4 Implementation Details
In our main experiments using the proposed DRS method, we set the temperature to 0.0 for all large language models to ensure reproducibility. Similarly, for the GPT-4o-mini evaluator, we set the temperature to 0.0 and top_p to 1.0 to maintain reliability and consistency in the model’s judgments.
For the two GPT series models, we use the official OpenAI API [^2] https://openai.com/ for inference. For the two open-source models, we use the weights released on HuggingFace [^3] https://huggingface.co/ and deploy them on a single NVIDIA RTX A6000 GPU using bfloat16 precision and utilize the generate function for text generation. (Additional details are provided in Appendix A).
@@ 5 GPT Evaluator
6 Experimental Results
To demonstrate the effectiveness of our proposed zero-shot DRS method, we test it on six datasets, totaling over 1,700 data points, using four different large language models. We conduct a fair and thorough comparison between our method and four baseline methods from previous research, with all accuracy scores presented in Table 3. Additionally, we adjust relevant parameters, such as temperature and the number of candidate questions, in the zero-shot DRS method and conduct multiple experiments. The results are shown in Figure 4 and Figure 3. Finally, considering GPU resources and API costs, we select a subset of test data for additional experiments with more models, as listed in Appendix C.
@@ 6 Experimental Results
Inference Time of DRS. In Table 4, we present the average runtime of GPT-3.5 and Gemma2-9B under different methods. For our DRS method, we include the runtime when the number of candidate questions is set to 2 or 3, as Figure 3 shows that optimal performance can already be achieved with 2 or 3 candidates. From Table 4, we observe that the DRS method, implemented based on a DFS search approach, results in longer runtimes compared to the few-shot CoT method, despite the use of effective pruning techniques. Specifically, when the number of candidate questions is 2, the runtime increases by 37% on GPT-3.5 and 25% on Gemma2-9B.
@@ 8 Conclusion
Acknowledgements
The authors would like to thank the anonymous reviewers for their discussion and feedback.
@@ Appendix C Additional Experiments
Appendix D DRS Prompts

### v5

@@ 1 Introduction
We conduct extensive experiments on six diverse datasets, comparing DRS with multiple baseline approaches across different types of questions and domains. The results demonstrate that DRS significantly improves reformulation accuracy across all tested LLMs. Additionally, we introduce a more reliable evaluation framework using GPT-4o-mini, replacing the previous Llama2-7B evaluator to ensure more accurate assessment of reformulation quality.
@@ 3.1 Entity Extraction and Filtering
Therefore, this process involves two following sub-steps:
@@
Table 1: The detail information of all experimental data for evaluating the performance on the question reformulation task. The Length is calculated based on the number of tokens after tokenization using Gemma2-9B.
@@ 3.2 DFS Combination Search and Structured Question Generation
In this step, our algorithm uses a DFS search approach to identify possible entity combinations that can be reformulated into new questions. During the generation process, we find that directly prompting the large language model to reshape the question often results in a low success rate due to the model’s potentially unlimited output content. To address this, we guide the large language model to generate structured outputs by restricting the entities required in the generated text. We ensure the reliability of the reformulated question by first generating a statement, then using that statement to create the question.
Additionally, to enhance the algorithm’s efficiency, we verify the entities immediately after generating the new question, preventing unnecessary subsequent steps. We also apply pruning to the DFS algorithm by limiting the total number of preselected questions and controlling the search depth.
@@ 4.1 Datasets
In this paper, we utilize the newly constructed high-quality dataset CouldAsk [^1] https://huggingface.co/datasets/wentingzhao/couldask, introduced by Zhao et al. (2024). This dataset comprises subsets from six diverse sources, including Yelp, BBC, SQuAD, Reddit, and others, featuring over 4,000 high-difficulty, high-quality data points filtered by both human annotators and GPT-4 OpenAI et al. (2024). For our study, we focus on over 1,700 unanswerable questions selected from all subsets as test data for model reformulation. The detailed information about our experimental data is provided in Table 1.
@@ 4.2 Large Language Models
To evaluate our methodology and compare it with baseline approaches from previous studies, we conduct all experiments using four different LLMs: GPT-3.5, GPT-4o-mini, Gemma2-9B and Qwen2.5-7B Team et al. (2024); Team (2024). The detailed information of different LLMs is shown in Appendix B.
@@ 4.3 Metric
(i) We use GPT-4o-mini as an evaluator to generate reasoning steps and assess whether the reformulated question can be answered using the relevant text. (ii) We use GPT-4o-mini as an entity detector to measure the overlap of entities between the reformulated question and the original question. If the number of overlapping entities exceeds half of the labeled entities in the dataset, we consider the reformulated question likely to help the inquirer obtain the desired information.
@@ 5 GPT Evaluator
6 Experiment Results
To demonstrate the effectiveness of our proposed zero-shot DRS method, we test it on six datasets, totaling over 1,700 data points, using four different large language models. We conduct a fair and thorough comparison between our method and four baseline methods from previous research, with all accuracy scores presented in Table 3. We incorporate the phrase "think step by step" to elicit the models’ chain-of-thought reasoning capabilities Kojima et al. (2023); Wei et al. (2023). Additionally, we adjust relevant parameters, such as temperature and the number of candidate questions, in the zero-shot DRS method and conduct multiple experiments. The results are shown in Figure 4 and Figure 3. Finally, considering GPU resources and API costs, we select a subset of test data for additional experiments with more models, as listed in Appendix C.
@@ 8 Conclusion
Acknowledgments
This research is supported by Optum Labs, DARPA ANSR program FA8750-23-2-0004, a National Science Foundation CAREER award #2339766, and University of California, Merced.
@@ Appendix A Implementation Details
In our main experiments using the proposed DRS method, we set the temperature to 0.0 for all large language models to ensure reproducibility. Similarly, for the GPT-4o-mini evaluator, we set the temperature to 0.0 and top_p to 1.0 to maintain reliability and consistency in the model’s judgments. For the two GPT series models, we use the official OpenAI API [^2] https://openai.com/ for inference.
For the two open-source models Gemma2-9B and Qwen2.5-7B, we use the weights released on HuggingFace [^3] https://huggingface.co/ and deploy them on a single NVIDIA RTX A6000 GPU using bfloat16 precision and utilize the generate function for text generation.
@@ Appendix B Large Language Models
Closed-Source Models. To balance experimental cost and effectiveness, we primarily use two widely adopted models from OpenAI: GPT-3.5 and GPT-4o-mini. In Appendix C, we select a subset for the experiment using GPT-4.
@@
Open-Source Models. We also perform experiments using two well-known models developed by Google and Alibaba: Gemma2-9B Team et al. (2024) and Qwen2.5-7B Yang et al. (2024); Team (2024). In Appendix C, we select a subset for the experiment using Llama3.1-70B and Gemma2-27B.
@@ Appendix C Additional Experiments
Appendix D Inference Time of DRS
In Table 5, we present the average runtime of GPT-3.5 and Gemma2-9B under different methods. For our DRS method, we include the runtime when the number of candidate questions is set to 2 or 3, as Figure 3 shows that optimal performance can already be achieved with 2 or 3 candidates. From Table 5, we observe that the DRS method, implemented based on a DFS search approach, results in longer runtimes compared to the few-shot CoT method, despite the use of effective pruning techniques. Specifically, when the number of candidate questions is 2, the runtime increases by 37% on GPT-3.5 and 25% on Gemma2-9B.
            Dataset             Human Evaluation             QA2             100 / 100 (100%)             BanditQA             100 / 100 (100%)             BBC             59 / 59 (100%)             Reddit             100 / 100 (100%)             Yelp             51 / 51 (100%)             SQuADv2             99 / 100 (99%)
Table 6: Human evaluation results on six datasets for assessing the meaningfulness and relevance of reformulated questions.
Appendix E Human Evaluation
In this paper, we propose a zero-shot DRS method to significantly improve the question reformulation task. However, a key concern arises: Do LLMs merely reformulate extremely simple questions that are unrelated to the passages but can be easily answered by the models themselves? For example, How to spell wasabi?
To thoroughly assess the effectiveness of the DRS method, we recruit three native English-speaking graduate students to evaluate all datasets. For subsets with over 100 samples, we randomly selected 100 reformulated questions for evaluation. A question is deemed valid if at least two students agree that it is meaningful and relevant to the passage. The human evaluation results are presented in Table 6.
The results clearly show that our proposed zero-shot DRS method generates reformulated questions that are nearly 100% meaningful and highly relevant to the given long context. Only one reformulated question within the SQuADv2 subset is deemed unmeaningful by three evaluators. This highlights the effectiveness of our method in faithfully preserving the user’s original intent and producing relevant reformulated questions.
Appendix F Experimental Prompts